OpenStudy (thebadyboy):

Susan is creating a solution. A formula for the solution calls for 7/10 mL of water. Susan wants to make 2/7 less of the solution than the formula creates. How much water should Susan use? Express your answer as a fraction in simplest form. [ ? ML ]

OpenStudy (yolo_king):

First find the greatest common denominator of 10 and 7

OpenStudy (yolo_king):

@Thebadyboy What is 10 * 7?

OpenStudy (thebadyboy):

70

OpenStudy (yolo_king):

Greatest common denominator is the common number and both 7 and 10 go into in this case it would be 70 right!

OpenStudy (yolo_king):

So now 7/10 is the same as 49/70 since all you have to do is multiply 7 by both sides and you get 49/70. Same with 2/7 multiply both sides by 10 and you will get?

OpenStudy (thebadyboy):

yes

OpenStudy (thebadyboy):

20/70

OpenStudy (yolo_king):

Exactly! So now grab both of the fractions and subtract them... So \[\frac{ 49 }{ 70 }-\frac{ 20 }{ 70 } =?\]

OpenStudy (yolo_king):

What do you get?

OpenStudy (yolo_king):

@Thebadyboy what is \[\frac{ 49 }{ 70 }-\frac{ 20 }{70 } = ?\]

OpenStudy (thebadyboy):

um 29/70

OpenStudy (yolo_king):

YES^^^^ That's the answer!!!!

OpenStudy (yolo_king):

and its already in simplest form.

OpenStudy (thebadyboy):

thx

OpenStudy (yolo_king):

No problem.

OpenStudy (anonymous):

I don't buy it. If a cookie recipe calls for a stick of butter to make a batch of cookies, then if I want to make half a batch of cookies, I would cut the amount of all the ingredients in half. So that stick of butter becomes 1/2 a stick of butter. Mathematically it would look like this: 1/2 * 1 = 1/2 Since she wants to reduce the amount of solution by 2/7, she needs 5/7 of the water called for in the formula.